# “It’s Always a Losing Game”: How Workers Understand and Resist Surveillance Technologies on the Job

### v1

@@ 3.1.3. Data Analysis
Using our filtered dataset, the first two authors manually checked for relevance to workplace surveillance. For example, we omitted posts that may talk about surveillance more generally but not in work-specific contexts or used keywords in a way that did not relate to surveillance (“computer monitor”). We also omitted posts that were deleted or removed. This narrowed our dataset further to 783 relevant posts.
From this dataset, we conducted a thematic analysis on 200 randomly selected posts and their respective comments using an inductive approach, first starting with our initial codes developed during our preliminary round of analysis and then using a codebook to conduct our full analysis. Higher-level codes captured the nature of posts (e.g. “Information Sharing, “Sharing of Personal Experience,” “Information Seeking”), the nature of surveillance (“Reason for Surveillance”, “Surveillance Type”, “Modality,” “Surveillant Actor”, ”Industry”) and Experiences of surveillance (“Sentiment towards Surveillance,” “Impact of Surveillance,” “Resistance Tactic”). Our initial Reddit analysis provided high-level insight into the various types of surveillance technologies in use, the range of industries implementing these technologies, workers’ understanding of what motivated managers to implement these technologies, and accounts of the impacts of monitoring and surveillance on workers. It also provided a broad view of the different resistance tactics workers employ and how they use Reddit to discuss the topic of workplace surveillance. To further protect worker anonymity, we did not include usernames or other identifying information and paraphrased or slightly altered all quotes included in this paper to reduce their searchability while still maintaining their meaning. Reddit posts are referenced by an author-assigned post ID (e.g., [R1]).
@@ 3.2.1. Participant Recruitment
To get a more nuanced and in-depth understanding of the experiences and perspectives of workers under surveillance, we paired our content analysis with in-depth interviews with workers and managers with first-hand experiences with workplace surveillance. We recruited participants by individually reaching out to post authors included in our dataset via Reddit’s private messaging system. We also posted a recruitment ad on 3 of the subreddits with moderator permission (/r/antiwork, /r/remotework, and /r/remotejobs). In the ad, we included a link to a screener survey to determine their eligibility for the study. The survey included questions about their Reddit username, country of residence, as well as their experiences and general opinions regarding workplace surveillance.
In total, we recruited 10 participants who came from a range of different work experiences, including in customer support, hospitality, AI data work, marketing, logistics, and construction. 6 were remote workers and three participants had management experience (P2, P4, P7). Participants came from the US, Canada, India, Mexico, Japan, and the Philippines. All had experiences with at least 1 type of workplace surveillance technology, including with audio and video monitoring, Emotion AI, digital activity tracking, and location tracking. To protect their anonymity, participants are referred to by their participant ID (e.g., [P1]). We provide detailed information about our participants in Table 2.
@@
P1 Customer Support Specialist (Remote) Finance US Audio and video monitoring, Digital Activity Tracking, Emotion AI
@@ 3.2.2. Data Collection and Analysis
Using the same codebook from our content analysis, we qualitatively coded our interviews in an iterative manner, adding two additional higher level codes including worker-driven alternatives (e.g. Opt Out, Transparency), and current trends of surveillance (e.g. Outsourcing, Lack of Disclosure). Our interviews extended the insights of our initial Reddit analysis, including capturing the experiences of surveilled international workers at US-based companies and managers who experienced both sides of surveillance, as the surveilled and the surveillant. The interviews also provided additional context on workers’ perceptions of data and privacy laws, as well as their thoughts on Reddit as a space for support.
@@ 3.3. Limitations
Lastly, we want to acknowledge the ethical issues around studying Reddit. While we adhered to the ethical standards of our institution’s IRB for our interview procedures, public Reddit data is not considered human subjects research according to US federal law and therefore not under the purview of our local IRB. Social media posters are typically not aware that their data can be used for research purposes without their explicit consent (Fiesler and Proferes, 2018). One way to mitigate this is to share the research output to the original community or poster. When we asked for permission to post our recruitment ad on subreddits, we let moderators know that we intended to share our findings with their community, and several moderators expressed interest in our project. However, reaching out to the original poster is more difficult. Due to the sensitivity of the issue of workplace surveillance, most posts are likely to be from “throwaway” accounts, intended to be used only temporarily before being abandoned (Ammari et al., 2019). In recruiting participants, we only reached out to a small group of individuals at a time, prioritizing the ones we felt were the most relevant to the purpose and scope of our research. We conducted our outreach with transparency and sensitivity to avoid undue pressure. Although we implemented strategies to protect posters’ anonymity through paraphrasing quotes and not disclosing usernames, a posters’ anonymity cannot be guaranteed (Bruckman, 2002; Markham, 2012).
@@ 4.1. Continuous and Opaque Surveillance Leads to Compounding Anxieties
While some workers on Reddit and in interviews believed that some monitoring can be beneficial for accountability and training purposes, many were concerned that the level of surveillance they experienced often went beyond what they believed was necessary for their jobs. Employees were often left in the dark concerning how these systems worked, what data they collected, and how this data was used which led to multiple anxieties about privacy and lack of trust in the workplace. Rather than improve performance and efficiency, there were multiple testimonials that described how WSTs actually hindered one’s productivity due to the immense stress they caused when these tools were used for continuous surveillance, control, and disciplinary action. Managers also expressed various anxieties about the addictive nature of surveillance and the lack of trust it created towards their workers. The compounding anxieties felt by both managers and workers contributed to a culture of distrust that ultimately harmed both.
@@ 4.1.1. Privacy Concerns and Lack of Transparency
Across Reddit and in our interviews, multiple remote workers who were being monitored with activity tracking software such as Hubstaff and ActivTrak reported having multiple privacy concerns, especially when these software continuously tracked their computer screens, mouse movements, keyboard activity, idle time, and app usage (ActivTrak, 2024; Hubstaff, [n. d.]). R9 called these systems ”invasive” due to the level of tracking they were subjected to, adding that the awareness of being constantly watched and having personal data continuously collected about them made them want to resign from their job. P3, a full-time senior-level design engineer based in India who worked as a remote-based data worker training AI chatbots on the side, was especially worried about the screenshot feature of Hubstaff. Knowing that his supervisor could view and take screenshots at any point of time, he was concerned about exposing his personal data on the platform:
@@
In their testimonials above, both P3 and P5 expressed numerous uncertainties about how they were being surveilled at work, indicating the lack of transparency and limited amount of information workers have about the nature of WSTs at their jobs. This seemed to be an explicit decision on the employers’ part. When asked whether she disclosed monitoring behavior to their workers, P7, a remote-based operations manager, said that she did not, as she believed that it would make workers feel more anxious. There also were managers that would make workers aware about some type of monitoring that was happening, but who would not offer the full picture. P2, who had experience working for the Canadian government as both an employee and manager, noted that sometimes the managers would notify workers that they were being monitored, but admitted that “that was maybe not followed all the time.”
P3 was told by his direct manager that Hubstaff would only track his website activity and take screenshots every 10 minutes and that she did not have access to the screenshots herself, but that they would be reviewed by an upper-level manager. However, P3 had suspicion that not only did his direct manager have access to these screenshots, but that Hubstaff was collecting more information than was initially disclosed. Through watching YouTube videos, he found out that HubStaff had the ability to collect data on his mouse movements, keyboard activity, idle time, and the apps he used. Additionally, he learned only through talking with his coworkers who had previously been disciplined that he must maintain a level of 85% productivity at all times or else risk being removed from the project. However, he was still unsure about how this score was calculated. While managers were reluctant to disclose information about surveillance to avoid making workers feel anxious, our testimonials show how the lack of transparency itself led workers to feel anxiety around the lack of privacy and not knowing the full extent of how these systems worked and how they were being used by their employers. This ultimately resulted in workers feeling more suspicious and distrustful of their employers.
@@ 4.1.2. Discipline and Punishment of Employees
Workers also reported instances when WSTs erroneously flagged them for certain behavior. P7, who previously worked as a remote-based English as a Second Language teacher, reported being subjected to emotion recognition software that would assess her bodily movements and facial expressions. Despite these systems being highly inaccurate, P7 described how getting flagged for “poor performance” by these systems had the additional impact of affecting her pay:
@@
P7 added that there was a three-strikes policy where they were subjected to disciplinary action, missed pay, and even the threat of firings for being flagged by the system multiple times. P1 had a similar experience with AI-based “Quality Assurance” systems that performed sentiment analysis on her calls and wrongly assessed her as being angry at customers based on her voice:
@@
In both cases, AI audio and video analyses flagged workers for poor performance based on assumptions that the workers felt were disconnected from reality. Their managers seemed to place more trust in the systems than in their employees and relied on the automated reports to determine whether more active surveillance was required or whether an employee should be let go. P2 stated that managers may even target specific workers more frequently if they have a personal vendetta against them: “If someone had an inkling that someone wasn’t doing something right or maybe even had a grudge, I think people were being monitored more than they should have been.
There were multiple reports of workers forgoing breaks to go to the bathroom or to get a drink of water due to the risk of not meeting productivity standards set by WSTs. Every time P5 needed to go to the bathroom or leave her computer even for a few minutes, she would pause the timer to make sure she was not “wasting the company[‘s] time.” However, she soon learned that she was caught between two predicaments: Take a break, and end up working longer hours, or avoid breaks all together to finish earlier but experience intense fatigue:
@@
“I do feel that I have to be productive every single minute. And that has affected my sleep schedule, that has affected my work-life balance in a way. And emotionally, it’s just draining. I feel forced in a way to be more productive than…my peers who have normal jobs…and all that while I’m isolated and working by myself at home […] In terms of psychological effects, I do think that it takes a toll on remote workers, because it makes you feel like you’re not off trust, it makes you feel isolated from your team. And it…makes you feel like you’re a robot that needs to be tracked every single time for everything they do as if your work is just a machine behind the computer.” — P5
Feelings of intense exhaustion were echoed by many workers on Reddit and throughout our interviews, and how overwork often led to other health issues, such as depression, anxiety, and panic attacks. Many ended up being pushed to a point where they were no longer able to work and were either let go or forced to quit their jobs. The few who were able to get disability accommodations were still harassed by their managers for taking required breaks. P2 stated that monitoring systems disproportionately impacted people with disabilities because they specifically target outliers outside of the standardized ”norms”: ”That would often result in invasive lines of questioning for people. Like, if someone had to go to the bathroom fairly regularly because they had IBS, or someone was dealing with a mental health crisis…you’re kind of attacking them for doing this that they need to do.” Further, when P2 requested a disability accommodation that would allow them to take more breaks, they were demoted and their contract did not get renewed:
You can absolutely receive accommodations from this organization, but as soon as you inconvenience them, since you’re on a short-term contract, you will just not be renewed. It’s happened to so many people, it happened to me eventually, I was demoted. Because I asked for an accommodation. I’d never had a poor work review. But since I asked for this accommodation, they put me down. They said, it’ll be easier for me. So I just put up to that…The laws are in place, but they haven’t structured to circumvent them in the first place.” — P2
@@ 4.1.3. Uneasy Relationships in the Workplace
WSTs created a culture of resentment and distrust amongst workers and managers when used as a disciplinary tool. On Reddit, workers reported having negative perceptions of their managers who heavily relied on WSTs, frequently describing them as ”micromanagers” [R6, R53, R77, P82]. P2 stated that the surveillance did not go away once they were promoted to a manager. Initially happy about getting promoted because they wanted “get out from under [the surveillance],” they soon learned that they still were being surveilled by their upper-level managers with same tracking software, although they did not have much insight into how they were being assessed.
They felt guilty about being in a position where they were doing the same surveillance they were previously subjected to. “Since it was so bureaucratic…you had to follow things a certain way, even if you had reservations about it.” Their promotion led them to face distrust from their former peers. They explained that even though their role technically was qualified to be part of the bargaining unit of their union, other union members would not trust them, even though they believed that their role could be an asset due to their “insider” role as a manager. Although P2 stated that they did not want to be a “cop” and avoided disciplining workers for infractions found through surveillance technologies, they still were required to report it to their higher-level managers. P1 echoed that her managers were also monitored to make sure they were enforcing surveillance policies despite their negative feelings about it. “Everything they did was surveilled too…just like us […] The trainers…you could tell the ones that didn’t agree with it because they were being forced to tell us to turn our cameras on.
As a manager, P4 noted that the convenience of viewing camera footage at his restaurant via his mobile app anytime and any place made him “addicted” to monitoring, which also made him more paranoid and distrustful of his workers. This anxiety and addiction on grew the few times he did catch workers doing what they were not supposed to do:
@@ 4.2.1. Commiseration
One of the most common and accessible tactics employed by workers who were facing surveillance was simply commiserating with other workers, both online or in person. P3, for example, said that he was in regular communication with some of his colleagues through video calls, where they would discuss strategies or other information, which are not privy to project managers. It was through talking to her coworkers that P7 learned tactics like mouse jiggling. Since activity-tracking apps typically measured productivity through tracking mouse movements and keyboard activity, some workers learned that they could bypass being flagged for inactivity by continuously moving their mousepad.
P5 similarly commiserated with her coworkers over a WhatsApp group. These platforms, which are separate from their company’s platforms, provided her and her coworkers with a safer space to talk about work issues. While she appreciated this support, P5 still felt isolated since many of her coworkers did not seem as concerned about the issue of surveillance: “I don’t think that they feel so strongly about this as I do. Because I voiced my concern many times, and they just keep saying, well, it is what it is, this is what we signed up for. [But] I don’t think that this is what we signed up for, you know? Although P5’s friends have suggested that she talk directly to her manager about intrusive monitoring, she feared that she would face disciplinary action if she does:
@@
Taking learnings from the advice she received from Reddit, P5 went back to her peers on WhatsApp to encourage them to take breaks when they needed to without worrying about the time tracker:“I think I’m an instigator. But I kind of told them, ‘guys don’t feel guilty if you have to stand up and go to the bathroom and take a little longer than usual because honestly, what’s five minutes that you are not clicking through anything?” Similarly, P9 commiserated with other coworkers in person about monitoring but faced a similar response. Those close to her were empathetic but told her to deal with it. Feeling isolated, she found the “antiwork” SubReddit because she wanted to find a community with people who had similar work experiences as her. She mostly browsed but also posted from time to time to show support and solidarity for other workers who are going through similar experiences: “Being able to feel supported on this kind of big issue that was affecting me did make me feel like I wasn’t alone […] it was really helpful to know, just that I wasn’t by myself.” Being a part of the antiwork community also changed how P9 navigated her job:
@@
For both P5 and P9, as well as many other workers, Reddit provided an accessible entry point for them to start questioning the presence of WSTs and discover tactics of resistance while receiving advice, resources, and support from other workers in anonymous way. This experience led both of them to face a shift in “ideology,” as P9 put it, that recognized the agency they had in shaping their working conditions and the wider sociopolitical aspects of their labor. Through commiseration, P5 was able to further transfer the knowledge she learned from Reddit to her coworkers on WhatsApp. By sharing experiences and strategies on these platforms, workers were able to foster a sense of community that helped mitigate feelings of isolation. This empowered individuals to start questioning and resisting against surveillance practices, while bringing others into the fold.

### v2

@@ 3.1.3. Data Analysis
Using our filtered dataset, the first two authors manually checked for relevance to workplace surveillance. For example, we omitted posts that discussed surveillance more generally but not in work-specific contexts or used keywords in a way that did not relate to surveillance (“computer monitor”). We also omitted posts that were deleted or removed. This narrowed our dataset further to 783 relevant posts.
From this dataset, we conducted a thematic analysis on 200 randomly selected posts and their respective comments using an inductive approach, first starting with our initial codes developed during our preliminary round of analysis and then using a codebook to conduct our full analysis. Higher-level codes captured the nature of posts (e.g. “Information Sharing,” “Sharing of Personal Experience,” “Information Seeking”), the nature of surveillance (“Reason for Surveillance,” “Surveillance Type,” “Modality,” “Surveillant Actor,” “Industry”) and Experiences of surveillance (“Sentiment towards Surveillance,” “Impact of Surveillance,” “Resistance Tactic”). Our initial Reddit analysis provided high-level insight into the various types of surveillance technologies in use, the range of industries implementing these technologies, workers’ understanding of what motivated managers to implement these technologies, and accounts of the impacts of monitoring and surveillance on workers. It also provided a broad view of the different resistance tactics workers employ and how they use Reddit to discuss the topic of workplace surveillance. To further protect worker anonymity, we did not include usernames or other identifying information and paraphrased or slightly altered all quotes included in this paper to reduce their searchability while still maintaining their meaning. Reddit posts are referenced by an author-assigned post ID (e.g., [R1]).
@@ 3.2.1. Participant Recruitment
To gain a more nuanced and in-depth understanding of the experiences and perspectives of workers under surveillance, we paired our content analysis with in-depth interviews with workers and managers with first-hand experiences with workplace surveillance. We recruited participants by individually reaching out to post authors included in our dataset via Reddit’s private messaging system. We also posted a recruitment ad on 3 of the subreddits with moderator permission (/r/antiwork, /r/remotework, and /r/remotejobs). In the ad, we included a link to a screener survey to determine their eligibility for the study. The survey included questions about their Reddit username, country of residence, as well as their experiences and general opinions regarding workplace surveillance.
In total, we recruited 10 participants who came from a range of different work experiences, including in customer support, hospitality, AI data work, marketing, logistics, and construction. 6 were remote workers and four participants had management experience (P2, P3, P4, P7). Participants came from the US, Canada, India, Mexico, Japan, and the Philippines. All had experiences with at least 1 type of workplace surveillance technology, including with audio and video monitoring, emotion AI, digital activity tracking, and location tracking. To protect their anonymity, participants are referred to by their participant ID (e.g., [P1]). We provide detailed information about our participants in Table 2.
@@
P1 Customer Support Specialist (Remote) Finance US Audio/Video Monitoring, Digital Activity Tracking, Emotion AI
@@ 3.2.2. Data Collection and Analysis
Using the same codebook from our content analysis, we qualitatively coded our interviews in an iterative manner, adding two additional higher level codes including worker-driven alternatives (e.g., “Opt Out,” “Transparency”), and current trends of surveillance (e.g., “Outsourcing,” “Lack of Disclosure”). Our interviews extended the insights of our initial Reddit analysis, including capturing the experiences of surveilled international workers at US-based companies and managers who experienced both sides of surveillance, as the surveilled and the surveillant. The interviews also provided additional context on workers’ perceptions of data and privacy laws, as well as their thoughts on Reddit as a space for support.
@@ 3.3. Limitations
Lastly, we want to acknowledge the ethical issues around studying Reddit. While we adhered to the ethical standards of our institution’s IRB for our interview procedures, public Reddit data is not considered human subjects research according to US federal law and therefore not under the purview of our local IRB (of Health and Services, [n. d.]). Social media posters are typically not aware that their data can be used for research purposes without their explicit consent (Fiesler and Proferes, 2018). One way to mitigate this is to share the research output to the original community or poster. When we asked for permission to post our recruitment ad on subreddits, we let moderators know that we intended to share our findings with their community, and several moderators expressed interest in our project. However, reaching out to the original poster is more difficult. Due to the sensitivity of the issue of workplace surveillance, most posts are likely to be from “throwaway” accounts, intended to be used only temporarily before being abandoned (Ammari et al., 2019). In recruiting participants, we only reached out to a small group of individuals at a time, prioritizing the ones we felt were the most relevant to the purpose and scope of our research. We conducted our outreach with transparency and sensitivity to avoid undue pressure. Although we implemented strategies to protect posters’ anonymity through paraphrasing quotes and not disclosing usernames, a posters’ anonymity cannot be guaranteed (Bruckman, 2002; Markham, 2012).
@@ 4.1. Continuous and Opaque Surveillance Leads to Compounding Anxieties
While some workers on Reddit and in interviews believed that some monitoring can be beneficial for accountability and training purposes, many were concerned that the level of surveillance they experienced often went beyond what they believed was necessary for their jobs. Employees were often left in the dark concerning how these systems worked, what data they collected, and how this data was used which led to multiple anxieties about privacy and lack of trust in the workplace. Rather than improve performance and efficiency, there were multiple testimonials that described how WSTs actually hindered workers’ productivity due to the immense stress they caused when these tools were used for continuous surveillance, control, and disciplinary action. Managers also expressed various anxieties about the addictive nature of surveillance and the lack of confidence it cultivated between themselves and their employees. The compounding anxieties felt by both managers and workers contributed to a culture of distrust that ultimately harmed both.
@@ 4.1.1. Privacy Concerns and Lack of Transparency
Across Reddit and in our interviews, multiple remote workers who were being monitored with activity tracking software such as Hubstaff and ActivTrak reported having multiple privacy concerns, especially when these software continuously tracked their computer screens, mouse movements, keyboard activity, idle time, and app usage (ActivTrak, 2024; Hubstaff, [n. d.]). R9 called these systems “invasive” due to the level of tracking they were subjected to, adding that the awareness of being constantly watched and having personal data continuously collected about them made them want to resign from their job. P3, a full-time senior-level design engineer based in India who worked as a remote-based data worker training AI chatbots on the side, was especially worried about the screenshot feature of Hubstaff. Knowing that his supervisor could view and take screenshots at any point of time, he was concerned about exposing his personal data on the platform:
@@
In their testimonials above, both P3 and P5 expressed numerous uncertainties about how they were being surveilled at work, indicating the lack of transparency and limited amount of information workers have about the nature of WSTs at their jobs. This seemed to be an explicit decision on the employers’ part. When asked whether she informed her workers about monitoring their behavior, P7 — a remote-based operations manager at a construction company who used GPS tracking to oversee employee use of company vehicles — explained that she did not disclose this information because she believed it would increase their anxiety. There also were managers that would make workers aware about some type of monitoring that was happening, but who would not offer the full picture. P2, who had experience working for the Canadian government as both an employee and manager, noted that sometimes the managers would notify workers that they were being monitored, but admitted that “that was maybe not followed all the time.”
P3 was told by his direct manager that Hubstaff would only track his website activity and take screenshots every 10 minutes and that she did not have access to the screenshots herself, but that they would be reviewed by an upper-level manager. However, P3 had suspicion that not only did his direct manager have access to these screenshots, but that Hubstaff was collecting more information than was initially disclosed. Through watching YouTube videos, he found out that HubStaff had the ability to collect data on his mouse movements, keyboard activity, idle time, and the apps he used. Additionally, he learned only through talking with his coworkers who had previously been disciplined that he must maintain a level of 85% productivity at all times or else risk being removed from the project. However, he was still unsure about how this score was calculated. While managers were reluctant to disclose information about surveillance to avoid making workers feel anxious, our testimonials show how the lack of transparency itself led workers to feel anxiety around their lack of privacy and not knowing the full extent of how these systems worked and how they were being used by their employers. This ultimately resulted in workers feeling more suspicious and distrustful of their employers.
@@ 4.1.2. Discipline and Punishment of Employees
Workers also reported instances when WSTs erroneously flagged them for certain behavior. P7, who previously worked as a remote-based English as a Second Language teacher, reported being subjected to emotion recognition software that would assess her bodily movements and facial expressions. Despite these systems being notoriously inaccurate, P7 described how getting flagged for “poor performance” by these systems had the additional impact of affecting her pay:
@@
P7 added that there was a three-strikes policy where they were subjected to disciplinary action, missed pay, and even the threat of termination for being flagged by the system multiple times. P1 had a similar experience with AI-based “Quality Assurance” systems that performed sentiment analysis on her calls and wrongly assessed her as being angry at customers based on her voice:
@@
In both cases, AI audio and video analyses flagged workers for poor performance based on assumptions that the workers felt were disconnected from reality. Their managers seemed to place more trust in the systems than in their employees and relied on the automated reports to determine whether more active surveillance was required or whether an employee should be let go. P2 stated that managers targeted specific workers more frequently when they have a personal vendetta against them: “If someone had an inkling that someone wasn’t doing something right or maybe even had a grudge, I think people were being monitored more than they should have been.”
There were multiple reports of workers forgoing breaks to go to the bathroom or to get a drink of water due to the risk of not meeting productivity standards set by WSTs. Every time P5 needed to go to the bathroom or leave her computer even for a few minutes, she would pause the timer to make sure she was not “wasting the company[’s] time.” However, she soon learned that she was caught in a predicament—take a break and end up working longer hours, or avoid breaks all together to finish earlier but experience intense fatigue:
@@
“I do feel that I have to be productive every single minute. And that has affected my sleep schedule, that has affected my work-life balance in a way. And emotionally, it’s just draining. I feel forced in a way to be more productive than…my peers who have normal jobs…and all that while I’m isolated and working by myself at home […] In terms of psychological effects, I do think that it takes a toll on remote workers, because it makes you feel like you’re not of trust, it makes you feel isolated from your team. And it…makes you feel like you’re a robot that needs to be tracked every single time for everything they do as if your work is just a machine behind the computer.” — P5
Feelings of intense exhaustion were echoed by many workers on Reddit and throughout our interviews, and some described how overwork led to other health issues, such as depression, anxiety, and panic attacks. Many ended up being pushed to a point where they were no longer able to work and were either let go or forced to quit their jobs. The few who were able to get disability accommodations were still harassed by their managers for taking required breaks. P2 stated that monitoring systems disproportionately impacted people with disabilities because they specifically target outliers outside of the standardized “norms”: “That would often result in invasive lines of questioning for people. Like, if someone had to go to the bathroom fairly regularly because they had IBS, or someone was dealing with a mental health crisis…you’re kind of attacking them for doing this thing that they need to do.” Further, when P2 requested a disability accommodation that would allow them to take more breaks, they were demoted and their contract did not get renewed:
You can absolutely receive accommodations from this organization, but as soon as you inconvenience them, since you’re on a short-term contract, you will just not be renewed. It’s happened to so many people, it happened to me eventually, I was demoted. Because I asked for an accommodation. I’d never had a poor work review. But since I asked for this accommodation, they put me down. They said, it’ll be easier for me. So I just put up to that…The laws are in place, but they have it structured to circumvent them in the first place.” — P2
@@ 4.1.3. Uneasy Relationships in the Workplace
WSTs created a culture of resentment and distrust amongst workers and managers when used as a disciplinary tool. On Reddit, workers reported having negative perceptions of their managers who heavily relied on WSTs, frequently describing them as “micromanagers” [R6, R53, R77, P82]. P2 stated that the surveillance did not go away once they were promoted to being a manager. Initially happy about getting promoted because they wanted “get out from under [the surveillance],” they soon learned that they still were being surveilled by their upper-level managers with the same tracking software, although they did not have much insight into how they were being assessed.
They felt guilty about being in a position where they were doing the same surveillance they were previously subjected to. “Since it was so bureaucratic…you had to follow things a certain way, even if you had reservations about it.” Their promotion led them to face distrust from their former peers. They explained that even though their role technically was qualified to be part of the bargaining unit of their union, other union members would not trust them, even though they believed that their role could be an asset due to their “insider” role as a manager. Although P2 stated that they did not want to be a “cop” and avoided disciplining workers for infractions found through surveillance technologies, they were still required to report it to their higher-level managers. P1 echoed that her managers were also monitored to make sure they were enforcing surveillance policies despite their negative feelings about it. “Everything they did was surveilled too…just like us […] The trainers…you could tell the ones that didn’t agree with it because they were being forced to tell us to turn our cameras on.”
As a restaurant manager, P4 noted that the convenience of viewing camera footage via his mobile app anytime and any place made him “addicted” to monitoring, which also made him more paranoid and distrustful of his workers. These feelings grew the few times he did catch workers doing something they were not supposed to:
@@ 4.2.1. Commiseration
One of the most common and accessible tactics employed by workers who were facing surveillance was simply commiserating with other workers, both online or in person. P3, for example, said that he was in regular communication with some of his colleagues through video calls, where they discussed strategies or other information, which project managers were not privy to. It was through talking to her coworkers that P7 learned tactics like mouse jiggling. Since activity-tracking apps typically measured productivity through tracking mouse movements and keyboard activity, some workers learned that they could bypass being flagged for inactivity by continuously moving their mousepad.
P5 similarly commiserated with her coworkers over a WhatsApp group. These platforms, which are separate from their company’s platforms, provided her and her coworkers with a safer space to talk about work issues. While she appreciated this support, P5 still felt isolated since many of her coworkers did not seem as concerned about the issue of surveillance: “I don’t think that they feel so strongly about this as I do. Because I voiced my concern many times, and they just keep saying, ‘Well, it is what it is, this is what we signed up for.’ [But] I don’t think that this is what we signed up for, you know?” Although P5’s friends have suggested that she talk directly to her manager about intrusive monitoring, she feared that she would face disciplinary action if she does:
@@
Taking learnings from the advice she received from Reddit, P5 went back to her peers on WhatsApp to encourage them to take breaks when they needed to without worrying about the time tracker: “I think I’m an instigator. But I kind of told them, ‘guys don’t feel guilty if you have to stand up and go to the bathroom and take a little longer than usual because honestly, what’s five minutes that you are not clicking through anything?” Similarly, P9 commiserated with other coworkers in person about monitoring but faced a similar response. Those close to her were empathetic but told her to deal with it. Feeling isolated, she found the “antiwork” subreddit because she wanted to find a community with people who had similar work experiences. She mostly browsed but also posted from time to time to show support and solidarity for other workers who are going through similar experiences: “Being able to feel supported on this kind of big issue that was affecting me did make me feel like I wasn’t alone […] it was really helpful to know, just that I wasn’t by myself.” Being a part of the antiwork community also changed how P9 navigated her job:
@@
For both P5 and P9, as well as many other workers, Reddit provided an accessible entry point for them to start questioning the presence of WSTs and discover tactics of resistance while receiving advice, resources, and support from other workers in an anonymous way. This experience led both of them to face a shift in “ideology,” as P9 put it, that recognized the agency they had in shaping their working conditions and the wider sociopolitical aspects of their labor. Through commiseration, P5 was able to further transfer the knowledge she learned from Reddit to her coworkers on WhatsApp. By sharing experiences and strategies on these platforms, workers were able to foster a sense of community that helped mitigate feelings of isolation. This empowered individuals to start questioning and resisting against surveillance practices, while bringing others into the fold.